# The KSBA moduli space of stable log Calabi–Yau surfaces

### v2

@@ 5.2. Mirror algebras
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. For every $p,q,r\in C(P)_{\mathbb{Z}}$ and $\beta\in NE(\mathcal{X}/\overline{\mathcal{X}})$, the structure constant $N_{pqr}^{\beta}(\mathcal{X})\in\mathbb{Q}$ is defined by
$N_{pqr}^{\beta}(\mathcal{X})=\begin{cases}\mathrm{deg}[\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau_{pqr}^{\beta},z)]^{\mathrm{virt}}&\quad\text{if}\,\,\,r\neq 0\\ 1&\quad\text{if}\,\,\,p=q=r=0\,\,\,\text{and}\,\,\,\beta=0\\ 0&\quad\text{otherwise.}\\ \end{cases}$
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. For every monoid ideal $I\subset NE(\mathcal{X}/\overline{\mathcal{X}})$ such that $NE(\mathcal{X}/\overline{\mathcal{X}})\setminus I$ is finite, the mirror algebra $R_{I}(\mathcal{X})$ is the free $S_{I}(\mathcal{X})=\mathbb{C}[NE(\mathcal{X}/\overline{\mathcal{X}})]/I$-module
@@
$\vartheta_{p}\cdot\vartheta_{q}=\sum_{r\in C(P)_{\mathbb{Z}}}\sum_{\beta\in NE(\mathcal{X}/\overline{\mathcal{X}})\setminus I}N_{pqr}^{\beta}(\mathcal{X})t^{\beta}\vartheta_{r}$ (5.5)
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. For every monoid ideal $I\subset NE(\mathcal{X}/\overline{\mathcal{X}})$ such that $NE(\mathcal{X}/\overline{\mathcal{X}})\setminus I$ is finite, the mirror algebra $R_{I}(\mathcal{X})$ is a commutative, associative $S_{I}(\mathcal{X})$-algebra with unit given by $\vartheta_{0}$.
@@
All the arguments in the proof of [48, Theorem 1.9] for projective degenerations $g:\mathcal{X}\rightarrow S$ as in §5.1 can be adapted to the set-up of open Kulikov degenerations to prove Theorem 5.5. Indeed, the proof of [48, Theorem 1.9] for a projective degeneration $g:\mathcal{X}\rightarrow S$ is based on the study of the moduli spaces $\mathcal{M}(\mathcal{X},\tau,z)$, where $\tau$ is a genus zero type of punctured maps with three marked points and an additional punctured point. For an open Kulikov degeneration $\pi:\mathcal{X}\rightarrow\Delta$, the corresponding moduli spaces $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau,z)$ are proper by Theorem 5.2 and so the non-compactness of $\mathcal{X}_{0}$ is not an issue. ∎
@@ 5.3. Structure constants from the canonical scattering diagram
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration, with tropicalization $(C(P),C(\mathscr{P}))$. By [49], the structure constants $N_{pqr}^{\beta}(\mathcal{X})$ can be computed in terms of broken lines in the canonical scattering diagram for $(\mathcal{X},\mathcal{X}_{0})$. The canonical scattering diagram $\mathfrak{D}_{\mathcal{X}}$ is a collection of walls in $C(P)$, which are cones $C(W)$ over line segments $W$ of rational slopes contained in 2-dimensional faces $\sigma_{W}$ of $\mathscr{P}$ in $P$. Every wall $C(W)$ is endowed with a kink $\kappa_{W}\in NE(\mathcal{X}/\overline{\mathcal{X}})$. If $W$ is contained in an internal edge $e$ of $\mathscr{P}$, then the kink $\kappa_{W}$ is equal to the class $[C_{e}]$ of the corresponding proper curve $C_{e}$ in the double locus of $\mathcal{X}_{0}$. Else, we have $\kappa_{W}=0$. Each wall $C(W)$ is also endowed with a power series $f_{W}\in\mathbb{Q}[\Lambda_{W}][\![NE(\mathcal{X}/\overline{\mathcal{X}})]\!]$, where $\Lambda_{W}$ is the 1-dimensional lattice of integral tangent vectors to $W$. The power series $f_{W}$ is a generating series of genus 00 punctured Gromov–Witten invariants $N_{\tau}$, indexed by wall types $\tau$ and virtually counting $\mathbb{A}^{1}$-curves,
@@
A broken line is a piecewise linear embedding $b:(-\infty,0]\rightarrow C(P)$, oriented by increasing values of the parameter, with finitely many domains of linearity $b_{j}$, each one contained in a 3-dimensional face of $C(\mathscr{P})$, and decorated with a monomial $c_{b_{j}}z^{m_{b_{j}}}$, where $c_{b_{j}}\in\mathbb{Q}[\![NE(\mathcal{X}/\overline{\mathcal{X}})]\!]$ and $m_{b_{j}}\in\Lambda_{b_{j}}$, where $\Lambda_{b_{j}}$ is the 1-dimensional lattice of integral tangent vectors to $b_{j}$, and with $-m_{b_{j}}$ pointing in the direction of the broken line. For the initial domain of linearity $b_{\mathrm{in}}$, we have $c_{b_{\mathrm{in}}}=1$ and $m_{b_{\mathrm{in}}}$ is called the initial direction of the broken line. Moreover, a bending from a domain of linearity $b_{j}$ to a domain of linearity $b_{j+1}$ occurs when the broken line crosses a wall $W$, and the monomial $c_{b_{j+1}}z^{m_{b_{j+1}}}$ is a monomial in the expansion of
@@
where $n_{W}$ is the primitive normal vector to $C(W)$, positive on the half-space containing $b_{j}$. Finally, we denote by $b_{\mathrm{out}}$ the final domain of linearity of $b$, and $c_{b_{\mathrm{out}}}z^{m_{b_{\mathrm{out}}}}$ the corresponding monomial. By [49, Theorem 6.1], for every $p,q,r\in C(P)_{\mathbb{Z}}$, we then have
@@
where the sum is over pairs of broken lines $b$, $b^{\prime}$ of initial directions $p$, $q$, ending at a common general point close to $r\in C(P)$, and with $m_{b_{\mathrm{out}}}+m_{b_{\mathrm{out}}^{\prime}}=r$. One can also work only on $P$ at height 1 in $C(P)$ by working with the radial projections on $P$ of the broken lines, called jagged paths
@@
Let $\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. Then, for every $p,q,r\in C(P)_{\mathbb{Z}}$ and $\beta\in NE(\mathcal{X}/\overline{\mathcal{X}})$, the structure constants $N_{pqr}^{\beta}$ only depends on the log central fiber $\mathcal{X}_{0}^{\dagger}$.
@@
As reviewed above, by [49, Theorem 6.1], $N_{pqr}^{\beta}(\mathcal{X})$ can be computed from the canonical scattering diagram, which only depends on $\mathcal{X}_{0}^{\dagger}$ by [49, Proposition 3.7]. ∎
@@
Lemma 5.6 does not immediately follows from Definition 5.3. Indeed, even if $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau,z)$ only depends on $\mathcal{X}_{0}^{\dagger}$, the definition of the relative obstruction theory, and so also the definition of the virtual fundamental class $[\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau,z)]^{\mathrm{virt}}$, depend a priori on $\mathcal{X}$.
@@ 6. Polarized mirrors of open Kulikov degenerations
After classifying in §6.1 the elementary flops between quasi-projective open Kulikov degenerations, we study in §6.2 how the structure constants $N_{pqr}^{\beta}(\mathcal{X})$ of the mirror algebra change under these flops. This is used in §6.3 to prove a key finiteness property: the mirror algebra of a quasi-projective open Kulikov degeneration $\pi:\mathcal{X}\rightarrow\Delta$ can be defined over $\mathbb{C}[NE(\mathcal{X}/\overline{\mathcal{X}})]$, and not just over the Artinian rings $\mathbb{C}[NE(\mathcal{X}/\overline{\mathcal{X}})]/I$. In §6.4-6.5, we define the secondary fan $\mathrm{Sec}(\mathcal{X}/\overline{\mathcal{X}})$ as a coarsening of the Mori fan $\mathrm{MF}(\mathcal{X}/\overline{\mathcal{X}})$. We show that there is a natural stacky fan $\mathcal{S}ec(\mathcal{X}/\overline{\mathcal{X}})$ with underlying fan $\mathrm{Sec}(\mathcal{X}/\overline{\mathcal{X}})$, and that the Proj of the mirror algebras of the various projective crepant resolutions of $\overline{\mathcal{X}}$ glue and extend to a polarized mirror family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{sec}},\mathcal{D}_{\mathcal{X}}^{\mathrm{sec}}+\epsilon\,\mathcal{C}_{\mathcal{X}}^{\mathrm{sec}})\rightarrow\mathcal{S}_{\mathcal{X}}^{\mathrm{sec}}$ over the toric Deligne–Mumford stack $\mathcal{S}_{\mathcal{X}}^{\mathrm{sec}}$ with fan $\mathrm{Sec}(\mathcal{X}/\overline{\mathcal{X}})$.
6.1. Elementary flops
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. By Proposition 3.36, $\pi:\mathcal{X}\rightarrow\overline{\mathcal{X}}$ is a Mori dream space, and so $\mathcal{X}$ admits finitely many small $\mathbb{Q}$-factorial modifications over $\overline{\mathcal{X}}$. Let $b:\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}$ be such a small $\mathbb{Q}$-factorial modification of $\mathcal{X}$ over $\overline{\mathcal{X}}$. Then, both $\mathrm{Nef}(\mathcal{X}/\overline{\mathcal{X}})$ and $\mathrm{Nef}(\mathcal{X}^{\prime}/\overline{\mathcal{X}})$ are maximal dimensional cones of the Mori fan $\mathrm{MF}(\mathcal{X}/\overline{\mathcal{X}})$. We say that $b$ is elementary if the intersection $\mathrm{Nef}(\mathcal{X}/\overline{\mathcal{X}})\cap\mathrm{Nef}(\mathcal{X}/\overline{\mathcal{X}^{\prime}})$ is a common codimension one face of $\mathrm{Nef}(\mathcal{X}/\overline{\mathcal{X}})$ and $\mathrm{Nef}(\mathcal{X}^{\prime}/\overline{\mathcal{X}})$, that is, if the face $NE(b)$ of $NE(\mathcal{X}/\overline{\mathcal{X}})$ spanned by the curves contracted by $b$ is an extremal ray. The following result classifies the possible elementary $\mathbb{Q}$-factorial modifications.
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration, and let $b:\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}$ be an elementary small $\mathbb{Q}$-factorial modification of $\mathcal{X}$ over $\overline{\mathcal{X}}$. Then, the exceptional locus of $b$ is a smooth rational curve $C\subset\mathcal{X}_{0}$, such that either
  • i)
    $C$ is an internal exceptional curve of an irreducible component of $\mathcal{X}_{0}$, or
  • ii)
    $C=V_{i}\cap V_{j}$ is an irreducible component of the double locus of $\mathcal{X}_{0}$, intersection of two irreducible components $V_{i}$, $V_{j}$ of $\mathcal{X}_{0}$, and such that $\deg N_{C|V_{i}}=\deg N_{C|V_{j}}=-1$.
In both cases, the normal bundle to $C$ in $\mathcal{X}$ is $N_{C|\mathcal{X}}=\mathcal{O}_{C}(-1)\oplus\mathcal{O}_{C}(-1)$, and $b$ is the standard $(-1,-1)$-flop of $C$. In particular, $\mathcal{X}^{\prime}\rightarrow\Delta$ is also a quasi-projective open Kulikov degeneration.

### v3

@@ 5.2. Mirror algebras
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. For every $p,q,r\in C(P)_{\mathbb{Z}}$, $\beta\in NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$, and $W$ as in (5.2), the structure constant $N_{pqr}^{\beta}(\mathcal{X})_{W}\in\mathbb{Q}$ is defined by
$N_{pqr}^{\beta}(\mathcal{X})_{W}=\mathrm{deg}[\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau_{pqr}^{\beta})_{W}]^{\mathrm{virt}}\,.$
As shown in [51, §7.2], the structure constants $N_{pqr}^{\beta}(\mathcal{X})_{W}$ are independent of the choice of $W$ as in (5.2). Consequently, we will omit $W$ from the notation and refer to the structure constants simply as $N_{pqr}^{\beta}(\mathcal{X})$.
Remark 5.4.
In the set-up of [51], for a degeneration $g:\mathcal{X}\rightarrow S$ as in §5.1, the structure constants $N_{pqr}^{\beta}(\mathcal{X})$ are initially defined in [51, Definition 3.21] using a choice of a point $z$ in the interior of the stratum $Z_{r}$. For $r=0$, we have $Z_{0}=\mathcal{X}$ in [51], and so the interior of $Z_{0}$ is the complement $\mathcal{X}\setminus\mathcal{X}_{0}$ of the central fiber $\mathcal{X}_{0}$. However, in our setting, where we only have a formal scheme $\mathcal{X}\rightarrow\Delta$, it is not meaningful to choose a point $z$ away from $\mathcal{X}_{0}$. As a result, the definition of the structure constants $N_{pqr}^{\beta}(\mathcal{X})$ given in [51, Definition 3.21] cannot be directly applied for $r=0$ in our context. Fortunately, a more general definition, using a choice of log point $W$ as in (5.2) is provided in [51, §7.2] (see also [52, Step I of the proof of Theorem 6.4]). This is actually the definition used in the proof of the associativity of the mirror algebra. Our definition of the structure constants $N_{pqr}^{\beta}(\mathcal{X})$ follows this more general approach, which can be applied even if $r=0$ since the point $z_{W}$ can then be chosen arbitrarily.
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. For every monoid ideal $I\subset NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ such that $NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})\setminus I$ is finite, the mirror algebra $R_{I}(\mathcal{X})$ is the free $S_{I}(\mathcal{X})=\mathbb{C}[NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})]/I$-module
@@
$\vartheta_{p}\cdot\vartheta_{q}=\sum_{r\in C(P)_{\mathbb{Z}}}\sum_{\beta\in NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})\setminus I}N_{pqr}^{\beta}(\mathcal{X})t^{\beta}\vartheta_{r}$ (5.5)
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. For every monoid ideal $I\subset NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ such that $NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})\setminus I$ is finite, the mirror algebra $R_{I}(\mathcal{X})$ is a commutative, associative $S_{I}(\mathcal{X})$-algebra with unit given by $\vartheta_{0}$.
@@
All the arguments in the proof of [51, Theorem 1.9] for projective degenerations $g:\mathcal{X}\rightarrow S$ as in §5.1 can be adapted to the set-up of open Kulikov degenerations to prove Theorem 5.6. Indeed, the proof of [51, Theorem 1.9] for a projective degeneration $g:\mathcal{X}\rightarrow S$ is based on the study of the moduli spaces $\mathcal{M}(\mathcal{X},\tau)_{W}$, where $\tau$ is a genus zero type of punctured maps with three marked points and an additional punctured point. For an open Kulikov degeneration $\pi:\mathcal{X}\rightarrow\Delta$, the corresponding moduli spaces $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau)_{W}$ are proper by Theorem 5.2 and so the non-compactness of $\mathcal{X}_{0}$ is not an issue. ∎
@@ 5.3. Structure constants from the canonical scattering diagram
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration, with tropicalization $(C(P),C(\mathscr{P}))$. By [52], the structure constants $N_{pqr}^{\beta}(\mathcal{X})$ can be computed in terms of broken lines in the canonical scattering diagram for $(\mathcal{X},\mathcal{X}_{0})$. The canonical scattering diagram $\mathfrak{D}_{\mathcal{X}}$ is a collection of walls in $C(P)$, which are cones $C(W)$ over line segments $W$ of rational slopes contained in 2-dimensional faces $\sigma_{W}$ of $\mathscr{P}$ in $P$. Every wall $C(W)$ is endowed with a kink $\kappa_{W}\in NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$. If $W$ is contained in an interior edge $e$ of $\mathscr{P}$, then the kink $\kappa_{W}$ is equal to the class $C_{e}]$ of the corresponding proper curve $C_{e}$ in the double locus of $\mathcal{X}_{0}$. Else, we have $\kappa_{W}=0$. Each wall $C(W)$ is also endowed with a power series $f_{W}\in\mathbb{Q}[\Lambda_{W}][\![NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})]\!]$, where $\Lambda_{W}$ is the 1-dimensional lattice of integral tangent vectors to $W$. The power series $f_{W}$ is a generating series of genus 0 punctured Gromov–Witten invariants $N_{\tau}$, indexed by wall types $\tau$ and virtually counting $\mathbb{A}^{1}$-curves,
@@
A broken line is a piecewise linear embedding $b:(-\infty,0]\rightarrow C(P)$, oriented by increasing values of the parameter, with finitely many domains of linearity $b_{j}$, each one contained in a 3-dimensional face of $C(\mathscr{P})$, and decorated with a monomial $c_{b_{j}}z^{m_{b_{j}}}$, where $c_{b_{j}}\in\mathbb{Q}[\![NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})]\!]$ and $m_{b_{j}}\in\Lambda_{b_{j}}$, where $\Lambda_{b_{j}}$ is the 1-dimensional lattice of integral tangent vectors to $b_{j}$, and with $m_{b_{j}}$ pointing in the direction of the broken line. For the initial domain of linearity $b_{\mathrm{in}}$, we have $c_{b_{\mathrm{in}}}=1$ and $m_{b_{\mathrm{in}}}$ is called the initial direction of the broken line. Moreover, a bending from a domain of linearity $b_{j}$ to a domain of linearity $b_{j+1}$ occurs when the broken line crosses a wall $W$, and the monomial $c_{b_{j+1}}z^{m_{b_{j+1}}}$ is a monomial in the expansion of
@@
where $n_{W}$ is the primitive normal vector to $C(W)$, positive on the half-space containing $b_{j}$. Finally, we denote by $b_{\mathrm{out}}$ the final domain of linearity of $b$, and $c_{b_{\mathrm{out}}}z^{m_{b_{\mathrm{out}}}}$ the corresponding monomial. By [52, Theorem 6.1], for every $p,q,r\in C(P)_{\mathbb{Z}}$, we then have
@@
where the sum is over pairs of broken lines $b$, $b^{\prime}$ of initial directions $p$, $q$, ending at a common general point close to $r\in C(P)$, and with $m_{b_{\mathrm{out}}}+m_{b_{\mathrm{out}}^{\prime}}=r$. Alternatively, one could consider only $P$ at height 1 in $C(P)$ by working with the radial projections on $P$ of the broken lines, called jagged paths [48, §4.5].
@@
Let $\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration, with tropicalization $(C(P),C(\mathscr{P}))$. Then, for every $p,q,r\in C(P)_{\mathbb{Z}}$ and $\beta\in NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$, the structure constants $N_{pqr}^{\beta}$ only depend on the log central fiber $\mathcal{X}_{0}^{\dagger}$.
@@
As reviewed above, by [52, Theorem 6.1], the structure constants $N_{pqr}^{\beta}(\mathcal{X})$ can be computed from the canonical scattering diagram, which depends only on $\mathcal{X}_{0}^{\dagger}$ by [52, Proposition 3.7]. ∎
@@
Lemma 5.7 does not immediately follow from Definition 5.3. Although $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau)_{W}$ depends only on $\mathcal{X}_{0}^{\dagger}$, the definition of the relative obstruction theory in [51, Definition 7.9], and consequently the definition of the virtual fundamental class $\mathcal{M}(\mathcal{X}_{0}^{\dagger},\tau)_{W}]^{\mathrm{virt}}$, depend a priori on $\mathcal{X}$.
@@ 6. Polarized mirrors of open Kulikov degenerations
After classifying in §6.1 the flops of relative Picard rank one between generic quasi-projective open Kulikov degenerations, we study in §6.2 how these flops affect the structure constants $N_{pqr}^{\beta}(\mathcal{X})$ of the mirror algebra. This is used in §6.3 to prove a key finiteness property: the mirror algebra of a quasi-projective open Kulikov degeneration $\pi:\mathcal{X}\rightarrow\Delta$ can be defined over $\mathbb{C}[NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})]$, and not just over the Artinian rings $\mathbb{C}[NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})]/I$. In §6.4-6.6, following [53], we introduce the secondary fan $\mathrm{Sec}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ as a coarsening of the Mori fan $\mathrm{MF}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$. We further establish the existence of a natural stacky fan $\mathcal{S}ec(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$, whose underlying fan is $\mathrm{Sec}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$. Finally, we show that the Proj of the mirror algebras of the various projective crepant resolutions of $\mathcal{X}^{\mathrm{can}}$ glue together, extending to a polarized mirror family $(\mathcal{Y}_{\mathcal{X}}^{\mathrm{sec}},\mathcal{D}_{\mathcal{X}}^{\mathrm{sec}}+\epsilon\,\mathcal{C}_{\mathcal{X}}^{\mathrm{sec}})\rightarrow\mathcal{S}_{\mathcal{X}}^{\mathrm{sec}}$ over the toric Deligne–Mumford stack $\mathcal{S}_{\mathcal{X}}^{\mathrm{sec}}$ with fan $\mathcal{S}ec(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$.
6.1. Flops of relative Picard rank one
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a generic quasi-projective open Kulikov degeneration. By Proposition 3.36, $\pi:\mathcal{X}\rightarrow\mathcal{X}^{\mathrm{can}}$ is a Mori dream space, and so $\mathcal{X}$ admits finitely many small $\mathbb{Q}$-factorial modifications over $\mathcal{X}^{\mathrm{can}}$. Let $b:\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}$ be such a small $\mathbb{Q}$-factorial modification of $\mathcal{X}$ over $\mathcal{X}^{\mathrm{can}}$. Since $K_{\mathcal{X}}=K_{\mathcal{X}^{\prime}}=0$, we will simply refer to such a $b$ as a flop. The nef cones $\mathrm{Nef}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ and $\mathrm{Nef}(\mathcal{X}^{\prime}/\mathcal{X}^{\mathrm{can}})$ are maximal dimensional cones of the Mori fan $\mathrm{MF}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$.
We say that $b:\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}$ is a flop of relative Picard rank one if the intersection $\mathrm{Nef}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})\cap\mathrm{Nef}(\mathcal{X}^{\prime}/\mathcal{X}^{\mathrm{can}})$ is a common codimension one face of $\mathrm{Nef}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ and $\mathrm{Nef}(\mathcal{X}^{\prime}/\mathcal{X}^{\mathrm{can}})$, that is, if the face $NE(b)$ of $NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ spanned by the curves contracted by $b$ is a ray. We classify the possible flops of relative Picard rank one in Theorem 6.7 below. To do this, we first review the elementary flops referred to as Type I and Type II in [91, pg. 12], and called M1 flops and M2 flops in [7, Definition 7.15]. We will follow the terminology of [7, Definition 7.15].
Definition 6.1.
Let $\mathcal{X}\rightarrow\Delta$ be an open Kulikov degeneration. An interior exceptional curve in $\mathcal{X}$ is a curve $C\simeq\mathbb{P}^{1}$, that is contained within a unique irreducible component $X^{v}$ of $\mathcal{X}_{0}$, and satisfies $C|_{X^{v}}^{2}=-1$.
Definition 6.2.
Let $\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. An M1 flop of $\mathcal{X}$ is a flop $b:\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}$ whose exceptional locus in $\mathcal{X}$ consists of a single interior exceptional curve. A multiple M1 flop is a flop $b:\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}$ whose exceptional locus in $\mathcal{X}$ consists of a finite disjoint union of interior exceptional curves.
Definition 6.3.
Let $\mathcal{X}\rightarrow\Delta$ be an open Kulikov degeneration. A (1,1)(-1,-1)-double curve in $\mathcal{X}$ is an irreducible component $C$ of the double locus of $\mathcal{X}_{0}$ such that $C\simeq\mathbb{P}^{1}$, and, denoting by $C_{1}$ and $C_{2}$ the irreducible components of the preimage of $C$ in the normalization of $\mathcal{X}_{0}$, and $X_{1}$, $X_{2}$ the irreducible components of the normalization of $\mathcal{X}_{0}$ containing $C_{1}$ and $C_{2}$ respectively, we have $C_{1}|_{X_{1}}^{2}=C_{2}|_{X_{2}}^{2}=-1$.
Definition 6.4.
Let $\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. An M2 flop of $\mathcal{X}$ is a flop $b:\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}$ whose exceptional locus in $\mathcal{X}$ consists of a single (1,1)(-1,-1)-double curve. A multiple M2 flop is a flop $b:\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}$ whose exceptional locus in $\mathcal{X}$ consists of a finite disjoint union of (1,1)(-1,-1)-double curves.
Remark 6.5.
M0 flops, whose exceptional locus is an internal (2)(-2)-curve, are also introduced in [91, pg. 12] and [7, Definition 7.15]. However, since this paper primarily focuses on generic open Kulikov degenerations, which do not contain internal (2)(-2)-curves, M0-flops do not occur in our setting.
The following result will be used to prove Theorem 6.7 on the exceptional locus of flops of relative Picard rank one.
@@
Let $\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration. Let $E$ be an interior exceptional curve and $C$ a (1,1)(-1,-1)-double curve. Then, the classes of $E$ and $C$ in $NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ are not proportional.
@@
By Lemma 3.23, the morphism $\mathcal{X}\rightarrow\mathcal{X}^{\mathrm{can}}$ is projective, implying that the classes $[E]$ and $[C]$ of $E$ and $C$ in $NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ are both non-zero. Consider the tropicalization $(P,\mathscr{P})$ of $\mathcal{X}_{0}$, and let $e$ be the edge of $\mathscr{P}$ corresponding to $C$, with adjacent vertices $v_{1}$ and $v_{2}$. Denote by $f_{3}$ and $f_{4}$ the faces of $\mathscr{P}$ that meet along $e$. Since $\mathscr{P}$ is a planar polyhedral decomposition, it is impossible for both $v_{1}=v_{2}$ and $f_{3}=f_{4}$ to hold simultaneously. As a consequence, there must be at least three components $X^{v}$ of $\mathcal{X}_{0}$ such that $X^{v}\cdot C\neq 0$. On the other hand, at most two components $X^{v}$ of $\mathcal{X}_{0}$ satisfy $X^{v}\cdot E\neq 0$, and so $[C]$ and $[E]$ cannot be proportional. ∎
Theorem 6.7.
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a generic quasi-projective open Kulikov degeneration, and let $b:\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}$ be a flop of $\mathcal{X}$ of relative Picard rank one. Then, either $b$ is a multiple M1 flop, or a multiple M2 flops. In particular, $\mathcal{X}^{\prime}\rightarrow\Delta$ is also a generic quasi-projective open Kulikov degeneration.
Proof.
Since $\mathcal{X}_{0}$ is generic, Lemma 3.32 ensures that no irreducible component of $\mathcal{X}_{0}$ contains an internal (2)(-2)-curve. Consequently, by the proof of [90, Lemma 3.2], the connected components of the exceptional locus of $b$ in $\mathcal{X}$ must be either interior exceptional curves or (1,1)(-1,-1)-double curves. Moreover, because $b$ is assumed to be of relative Picard rank one, the classes of all these curves are proportional. However, by Lemma 6.6, the curves classes of an interior exceptional curve and of a (1,1)(-1,-1)-double curve cannot be proportional. It follows that the connected components of the exceptional locus of $b$ consist either entirely of interior exceptional curves, or entirely of (1,1)(-1,-1)-double curves. Finally, the genericity of $\mathcal{X}^{\prime}\rightarrow\Delta$ follows from the local calculations in [7, §7C], which imply that the periods of open Kulikov surfaces remain invariant under both M1 and M2 flops. ∎
We show below that if the irreducible components $X^{v}$ of $\mathcal{X}_{0}$ are normal, and the double intersections $X^{v}\cap X^{v^{\prime}}$ are connected, then any flop of relative Picard rank one must be either an M1 flop or an M2 flop. In general, it may not always be possible to factorize a multiple M1 flop or a multiple M2 flop into a sequence of M1 and M2 flops defined as above within the quasiprojective category, though it is always possible analytically.
Theorem 6.8.
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a generic quasi-projective open Kulikov degeneration. Assume that every irreducible component $X^{v}$ of the central fiber $\mathcal{X}_{0}$ is normal, and that all the double intersections $X^{v}\cap X^{v^{\prime}}$ are connected. Then, every flop $b:\mathcal{X}\dashrightarrow\mathcal{X}^{\prime}$ of relative Picard rank one is either an M1 flop or an M2 flop.
Proof.
By Theorem 6.7, it suffices to show that two disjoint interior exceptional curves or two disjoint (1,1)(-1,-1)-double curves cannot have positively proportional curve classes.
Let us first consider the case of two disjoint (1,1)(-1,-1)-double curves $C_{1}$ and $C_{2}$. As the irreducible components of $\mathcal{X}_{0}$ are normal and the double intersections are connected, there exists $v$ and $w\neq v$ such that $C_{1}=X^{v}\cap X^{w}$. Moreover, $X^{v}$ and $X^{w}$ are exactly the irreducible components of $\mathcal{X}_{0}$ that have negative intersection with $C_{1}$. If $C_{2}$ is positively proportional to $C_{1}$, then $C_{2}$ also intersects negatively $X^{v}$ and $X^{w}$, and so $C_{2}=X^{v}\cap X^{w}=C_{1}$, contradicting that $C_{1}$ and $C_{2}$ are disjoint.